### v2

# Dynamically-decoupled hyper-Ramsey spectroscopy of optical clock transitions

Abstract
Hyper-Ramsey protocols have been successfully implemented on ultra-narrow optical clock transitions to reduce systematic frequency-shifts induced by AC-Stark shift and amplitude pulse variation. However, the compensation remains imperfect against laser probe intensity fluctuation, decoherence and unsuited for external quasi-static or low frequency noise perturbations. Here, we address these limitations by employing dynamical-decoupling methods composed by multiple rotary Hahn-echo pulses toggling probe frequency detunings between opposite signs during interrogating laser pulses. Time-optimized Uhrig sequences of refocusing pulses produce highly contrasted and robust hyper-Ramsey interferences against low-frequency noise distortions caused by environmental factors and imperfections in the probe parameters. Dynamically-decoupled SU(2) hyper-clocks pave the way to universal noise-resilient quantum sensors, unveiling fault-tolerant quantum metrology to track fundamental symmetries and search for new physics beyond the Standard Model.
The design of robust optical qubits that are resistant to environmental noise and imperfections by engineering light-matter interaction is expected to lead to new frontiers for quantum simulation and computation, time and frequency metrology. In parallel, fault-tolerant quantum sensors will unlock new perspectives for exploring fundamental symmetries and the search for new physics beyond the Standard Model [1, 2, 3, 4]. Laser metrology, which is based on ultra-narrow optical clock transitions, is a prime example of a highly active research field that is currently facing a new and challenging task [5, 6]. This involves not only the evaluation of frequency-shifts in optical lattices or in single ion clocks with a fractional inaccuracy at the relative level of [7, 8, 9, 10] but also to realize highly stable clock lasers with suppressed sensitivity to phase and intensity laser noise [11, 12, 13]. In the ongoing pursuit to enhance metrological performances of optical clocks, highly charged ions (HCI) are emerging as promising candidates to achieve a fractional frequency uncertainty below . Their high-order multipolar electronic transitions have much less sensitivity to external perturbations induced by black-body radiation and external electromagnetic fields [14, 15, 16]. Furthermore, spectroscopic performances can be enhanced through the optimization of control methods based on periodic driving, composite pulses and dynamical decoupling techniques used in quantum information processing [17, 18, 19, 20, 21, 22]. These techniques provide efficient quantum engineering solutions to reduce specific spin-spin interactions and distortions from many-body effects in experimental atomic physic platforms [23, 24, 25].
In the field of laser spectroscopy, various methods have been explored to shield a qubit-clock transition from external influences. Among these, hyper-Ramsey (HR) spectroscopy denoted as HR3π, was proposed in 2010 [26] and demonstrated experimentally in 2012 [27]. This method drastically reduces AC-Stark shifts induced by laser coupling to off-resonant atomic states and neighboring Zeeman sub-levels attributed to a small elliptic polarization of the probe [27, 28]. A modified version of HR spectroscopy was proposed and demonstrated in 2016 within optical lattices probing ultra-narrow clock transitions of bosons eliminating probe-field-induced AC-Stark shifts by three orders of magnitude even with significant errors in shift compensation [29]. Single-ion optical clocks based on 171Yb+ [30], 40Ca+ [31] and 176Lu+ [32] have exploited HR spectroscopy reporting an impressive low uncertainty budget at the relative level of accuracy. However, the asymmetric temporal position of the intermediate refocusing pulse in HR spectroscopy is required for the clock interferometer to be sensitive to the detuning of the probe beam with the highest signal contrast. It is opposed to the Mach-Zehnder atomic interferometer which eliminates not only inhomogeneous broadening effects but sensitivity to frequency offsets [33]. Consequently, the original HR scheme still exhibits a fundamental technical drawback suffering from residual probe frequency offsets and pulse errors associated to weak decoherence [34] requiring a combination of several sequences of phase-shifted composite pulses to eliminate some of these distortions [35, 36, 37]. Robustness of HR spectroscopy and hybrid schemes were finally investigated in presence of trapped-ion heating processes [38] and in an optically dense medium [39].
In this letter, we present a significant improvement of HR spectroscopy by inserting a single phase-shifted refocusing pulse to generate dynamical decoupling through symmetry considerations [40, 41, 42]. The symmetrization of the temporal position of a single refocusing pulse between Ramsey pulses generates a dynamical-decoupling effect which suppresses environmental low frequency noise impact and decoherence-induced frequency-shifts on interferences. Our new interrogation scheme denoted as dynamically-decoupled HR spectroscopy (DDHR3π), presented in Fig. 1, is merging composite pulse laser spectroscopy [26, 37, 43, 44, 45] with spin-echoes [46, 47] to preserve robustness against probe-induced light-shifts and frequency offsets and drifts. We introduce an intermediate rotary Hahn-echo pulse [48], coupled to a negative detuning [49, 50, 51, 52], which suppresses low noise fluctuation and frequency offsets that alter the optical qubit excitation probability by approximately the same amount in both free evolution zones. At the same time, residual clock frequency-shifts related to decoherence and laser-probe-intensity fluctuations that can not be compensated by the original HR protocol [34, 53] are suppressed. Our DDHR scheme is then extended to multiple refocusing pulses restoring sensitivity of quantum interferences to the laser probe frequency detuning at the output of the interferometer as demonstrated in refs [54, 55].
The original hyper-Ramsey HR3π protocol is implemented by using a composite pulse sequence without sign modification of the laser probe frequency detuning as reported in Fig. 1(a)-(b). Our new DDHR3π protocol is realized by inserting an intermediate refocusing pulse with a negative laser probe detuning (i.e the laser probe is red-detuned from the perturbed qubit-clock resonance) applied midway between Ramsey pulses as . Composite pulses and associated laser probe parameters including distortions are reported in Fig. 1(a)-(c). Swapping the probe detuning sign is realized by the intermediate time-reversal pulse action on the quantum system evolution. The laser probe frequency detuning can suffer from uncompensated residual light-shifts associated to potential static frequency offsets or residual frequency drifts due to technical defects as described in Fig. 1. The residual light-shift is induced by excitation pulses while a technical defect as a frequency offset or drift is always present over the entire sequence of pulses. The propagator matrix of the -th pulse (l) driving SU(2) spinor dynamics reads [44, 45]
with a pulse phase related to the Rabi frequency and where means we can apply a positive or a negative effective probe detuning where . Cayley-Klein phase angles are introduced as:
The pulse area is with a generalized Rabi frequency denoted as . The free evolution propagator while switching laser fields is
Transition probabilities associated to our protocols are computed by multiplying propagator matrices leading to:
including a phase-shifted refocusing pulse with . Exact analytic solutions of transition probabilities can be found in [45]. We plot HR3π and DDHR3π interference fringes versus the probe frequency detuning in Fig. 2(a)-(b) and interference contrast of HR3π and DDHR3π signals in Fig. 3 based on Eq. (4) and Eq. (5). The DDHR3π protocol is robust to a large pulse area variation while the HR3π scheme has already lost half of its population fraction. This is the first evidence of a robust sequence of composite pulses realized by a temporal symmetrization of laser excitation pulses on the optical qubit-clock transition.
We plot in Fig. 4 the clock frequency-shift as a function of a residual probe-induced light-shift including or not a small correction by a decoherence term . In Fig. 4, the HR3π frequency-shift of the central fringe exhibits a nonlinear cubic dependence with the probe-induced light-shift [26, 27] while the DDHR3π frequency-shift is immune. This is the second evidence of a robust sequence of composite pulses by inserting a modulation of the laser probe frequency detunings with opposite signs into a sequence of composite pulses [52, 51].
Also, we plot the central fringe frequency-shift with respect to residual probe-induced light-shift using Optical-Bloch equations including a decoherence term describing a finite laser probe linewidth [34, 36]. We report our results in Fig. 4 demonstrating that decoherence does not affect the reliability of the DDHR3π protocol in the presence of residual light-shifts, while the HR3π protocol breaks down in the full suppression of probe-induced light-shifts [34].
Then, we plot the interference contrast and sensitivity of HR3π and DD-HR3π protocols to a small frequency offset or drift related to technical defects of the laser probe or external distortions during the interrogation process. In Fig. 5(a), the HR3π protocol exhibits a steepest linear slope of the frequency-shift versus a residual probe detuning offset. The DDHR3π protocol unveils a non linear sensitivity to probe frequency detuning offsets even in presence of variation of pulse areas by several percents. Furthermore, robust dispersive DDHR error signals (DDDHR3±φ) can be produced by substracting two or more phase-shifted transition probabilities [35, 36, 37, 45]. A DDDHR3±π/4 error signal is reported in Fig. 5(a) removing the residual nonlinear dependence with the probe-induced frequency offset. A drastic improvement of interference contrast by the DDHR3π scheme versus the free evolution time T is presented in Fig. 5(b).
We now turn our attention to the influence of probe-laser-intensity fluctuations on HR3π and DDHR3π spectroscopy. Following a model from [53], we have investigated the effect of the fluctuation of the Rabi probe field coupled to probe-induced light-shifts on these composite pulse protocols and plotted the results in Fig. 6. The DDHR3π protocol is still immune with an error-free clock operation (solid red stars) while the HR3π protocol (blue solid dots) is compromised with a clock frequency-shift acquiring a lower-order quartic dependence with Rabi field fluctuations. This is another demonstration of fault-tolerance of our DDHR composite pulse protocol.
We finally propose to extend noise resilient DDHR spectroscopy to sequences with multiple refocusing pulses inspired by NMR techniques from periodic driving control [20, 22] to dynamical-decoupling [56, 57, 58, 59, 60, 61, 62, 63].
Dynamical decoupling methods aim not only to suppress dephasing errors from the environment but also to eliminate undesired quasi-static or time-dependent interactions by employing control sequences of multiple Hahn-echo refocusing pulses on atomic systems like qubits [64, 65, 66, 67, 68] and matter-waves [69]. Usually, a single Hahn-echo is sufficient in the case of quasi-static interactions (see Fig. 5(b)). However, low noise fluctuation can cause external interactions to vary [70]. Consequently, it is imperative to apply rotary Hahn-echoes repeatedly, preferably at a rate that is fast in comparison to the fluctuation time scale.
To simulate dynamical decoupling sequences of pulses in presence of low noise perturbation, we establish first a set of expansion rules for symmetrization of the qubit-clock dynamics driven by phase-shifted Hahn refocusing pulses toggling signs of frequency-shifted probe detunings as following:
- An odd number of phase-shifted Hahn-refocusing pulses encapsulated by free evolution zones with opposite sign can be inserted into Eq. 5 as
- An equal number of positive and negative free evolution zones symmetrically positioned over the entire sequence of pulses.
- An equal number of pulse areas made of positive and negative probe frequency detunings.
Then, in our simulations, we introduce different noise sources as a large imperfection of the Rabi field coupling and AC-Stark shift while inserting a variable frequency offset . We treat these noise sources as quasi-static, where each noise is modeled by a random value from a Gaussian distribution with standard deviation.
Following our expansion rules, DDHR spectroscopy incorporate quantum control protocols related to conventional dynamical decoupling like periodic-DD (PDDHR), concatenated-DD (CDDHR) interrogation protocols and Uhrig-DD (UDDHR) time-optimized sequences of rotary Hahn-echo pulses. They effectively disconnect laser probe and decoherence inducing frequency-shifts of interferences in agreement with the results of Fig. 4, Fig. 5 and Fig. 6.
We also introduce quantum logic schemes synthesizing noise filters based on Walsh-Hadamard patterns [71, 72].
The generation of Walsh-Hadamard-Hahn refocusing building-blocks (where indicates sequential, Hadamard and Paley matrix row ordering) is realized by a recursive application (initialization stage ) such that:
Following Eq. 6, we generate sequences of pulses in a sequential order shown in the upper part of Fig. 7. We have identified some pulse sequences of (k) equivalent to PDDHR protocols including 3 and 7 rotary Hahn-echo pulses or to CDDHR sequences including respectively 5 and 21 pulses.
Particular pulse sequences of WAHAHA and CDD protocols are formally equivalent to aperiodic self-similar Thue-Morse (TM) sequences [73]. The TM algorithm was applied for the first time to probe a Rabi clock resonance of a single trapped 171Yb+ ion minimizing low frequency noise fluctuation of the optical reference cavity [74].
The lower part of Fig. 7 shows interference contrast of HR3π, PDDHR3π and UDDHR3π protocols versus an increasing number of refocusing pulses. Replacing the PDDHR3π scheme by a UDDHR3π sequence of N refocusing pulses within magic time intervals following T (for ) produces the highest interference contrast against noise.
DDHR spectroscopy is a fault-tolerant interrogation protocol that has been designed to outperform hyper-Ramsey robustness against residual probe-induced frequency-shifts in presence of probe intensity fluctuation, residual drifts and noisy electromagnetic trapping fields. Optimal quantum control methods may be used for error-robust DDHR spectroscopy tailored to specific noise sources [75, 76]. Near-optimal DDHR schemes may be explored by nesting DD sequences for contrast optimization [77] while erasing residual technical pulse distortions by autobalanced DDHR spectroscopy [45]. Executing efficient spin squeezing in optical clocks may require multiple refocusing pulses and compensation of errors [78, 79, 80, 81] that might be compatible with our DDHR protocols.
Optical tweezer clocks [82, 83], Coulomb ion crystals [84, 85, 86], multiple highly-charged ion clocks [14, 15, 87], 229Th3+ single-ion nuclear clocks [88, 89] and neutral atom lattice clocks based on optical high-order multipolar transitions [90, 91] should benefit from interrogation schemes preserving long-lived atomic interferences against inhomogeneities and broadening mechanisms. DDHR spectroscopy may finally enable robust quantum sensing and metrology with ultracold atoms and molecules offering higher sensitivity to track dark matter and test new physics beyond the Standard Model [4, 92, 93].
-Acknowledgement- T. Z-W thanks J. Ye and J.A. Jones for discussions. D. W. thanks CQT/MoE, Grant No. R-710-002-016-271, for financial support.